(1 point) Find the point on the line 4x + 7y + 3 = 0 which is closest to the point (4,-1) Answer is
Answer #1

om Id 4 = 42+74 +3=0 closest distance on ho line the perpendicular distance from the pont the need to find P. to to 4. that AH Ap has slope Passes through (4,-1) and ş [y- (-1)) = 7/7 (2-4) yt! 7 (x-4) Eau 9 Yyty - 2x - 28 7. - Yy -32=0 L2 4 & ₂ ge